Question 11 Brad Swifty Corporation sells two types of computers; one is designed for audio applications and the other for video applications. Swifty incurs $348250 in fixed costs Per-unit data on the two products is presented blow: Unit data Selling price Variable costs Contribution margin Sales mix Audio computer $1530 1050 $480 75% Video computer $1770 1220 $550 25% What will be the total contribution margin at the break-even point? $238800 $27362:5 $348250 $512425
Answer #1

Answer : $348,250

At Break even point always contribution = Fixed cost

Given that fixed cost = $348,250

So Contribution = $348,250

Selling price per unit = 1530 + 1770 = 3300

Variable cost = 1050 + 1220 = 2270

Contribution = 480 + 550 = 1030

1 - 1030

? - 348,250

348,250 / 1030 * 1

338.11 units

Sales = 338.1 1 * 3300 = 1,115,763

Contribution margin = contribution / sales *100

= 1030 / 3300

Contribution margin = 31.212%

Contribution margin = 1,115,763 * 31.212 = 348,251 (approx)